OpenStudy (anonymous):

write the equation for a parabola of which has a vertex of (1, –3)?

OpenStudy (turingtest):

the equation of a parabola with vertex\[(h,k)\]is\[y-k=a(x-h)^2\]use that to make an equation.

OpenStudy (anonymous):

i know tht part but i dont know the number to put for a

OpenStudy (turingtest):

anything you want! it doesn't change the vertex :)

myininaya (myininaya):

it doesn't matter you just need to write an equation for a parabola

myininaya (myininaya):

oh lol

OpenStudy (anonymous):

oh ok

OpenStudy (turingtest):

you could turn it sideways too, if you wanted. still wouldn't change the vertex

OpenStudy (turingtest):

but you'd have to mess with the variables a little.

OpenStudy (anonymous):

y = a(x+3) + 1 where a can be anything except 0.

myininaya (myininaya):

right gj meverette because then it wouldn't be a parabola lol